# Latent-space variational data assimilation in two-dimensional turbulence

### v1

@@ Abstract
Starting from limited measurements of a turbulent flow, data assimilation (DA) attempts to estimate all the spatio-temporal scales of motion. Success is dependent on whether the system is observable from the measurements, or how much of the initial turbulent field is encoded in the available measurements. Adjoint-variational DA minimises the discrepancy between the true and estimated measurements by optimising the initial velocity or vorticity field (the ‘state space’). Here we propose to instead optimise in a lower-dimensional latent space which is learned by implicit rank minimising autoencoders. Assimilating in latent space, rather than state space, redefines the observability of the measurements and identifies the physically meaningful perturbation directions which matter most for accurate prediction of the flow evolution. When observing coarse-grained measurements of two-dimensional Kolmogorov flow at moderate Reynolds numbers, the proposed latent-space DA approach estimates the full turbulent state with a relative error improvement of two orders of magnitude over the standard state-space DA approach. The small scales of the estimated turbulent field are predicted more faithfully with latent-space DA, greatly reducing erroneous small-scale velocities typically introduced by state-space DA. These findings demonstrate that the observability of the system from available data can be greatly improved when turbulent measurements are assimilated in the right space, or coordinates.
1 Introduction
Assimilating experimental data into numerical simulations improves the fidelity of the simulations and enables nonintrusive access to all the scales of the estimated flow. However, the estimation of turbulence from limited measurements is a difficult ill-posed problem (zaki2025). Turbulence presents challenges such as the chaotic nature of the forward and dual problems, the non-uniqueness of solutions consistent with the measurements, and the introduction of erroneous small-scale velocities that decay over the solution trajectory.
@@ 1 Introduction
The term latent space refers to a low-dimensional and interpretable representation of the turbulent field, whether by familiar modal decompositions (taira2017modal), or by non-linear autoencoder transformations (brunton2020). Physical insights of complex dynamical systems can be gleaned from these low-dimensional latent spaces (Fukami2023). Rank-minimising autoencoders have successfully learned parsimonious representations of chaotic systems such as the Lorenz system, the Kuramoto-Sivashinsky equation and the lambda-omega reaction-diffusion system (Zeng_2024). The latent spaces of such autoencoders have yielded insights on the nature of bursting events and the dynamical relevance of unstable periodic orbits in forced two-dimensional (2D) turbulence (cleary2025).
@@
In the context of turbulence estimation, data-driven methods have been used to directly super-resolve limited instantaneous observations to the full flow state (e.g. Fukami_Fukagata_Taira_2019). Another super-resolution study considered a time-history of scarce measurements and inferred the pressure field of forced isotropic turbulent flow (williams2024). Machine-learning and classical DA methods to estimate turbulence have been pursued mostly separately, with a few exceptions including the following examples: DU2023 compared the estimation of wall turbulence using physics-informed neural networks and adjoint-variational techniques. Page_2025 modified the training of super-resolution networks to incorporate a time-forward Navier-Stokes evolution in the output, and a comparison to future data. Most recently, (weyrauch2025) used the output of this super-resolution network as an initial guess to adjoint-variational DA. While this last effort has interfaced the data-driven methods and adjoint-variational DA, the two techniques were not fully integrated.
@@
In the present work, we integrate the two turbulence estimation strategies. We exploit a learned space that is discovered using data-driven methods and the physics constraints of adjoint-variational DA, to significantly enhance the observability of turbulence systems and their estimation across a wider range of scales. In §2, we outline the flow configuration studied and summarise the standard variational DA procedure. In §3, we present the proposed latent-space DA procedure. In §4, we compare and interpret the improved performance of latent-space DA. We conclude in §5.
2 State-space data assimilation
@@
We consider Kolmogorov flow, which is monochromatically forced 2D turbulence on a square and doubly periodic domain (Chandler2013). The out-of-plane vorticity $\omega=\partial_{x}v-\partial_{y}u$ defines the flow state in state space and evolves according to
[…]
@@
Figure 1: Schematic of latent-space data assimilation. The latent representation $\eta$ is mapped to state space $\omega$ by the decoder $\mathcal{F}_{\mathcal{D}}$, where the adjoint field $\omega^{\dagger}$ is computed. The grey arrow closes the standard, state-space ‘loop’. In latent-space assimilation, the latent state is updated by the transformed adjoint field $\eta^{\dagger}$. Sensor resolution is indicated by the lattice of black dots in the lower left corner of $\omega$.
@@
Our objective is to estimate the flow state $\omega^{*}_{0}$ which, when evolved using the Navier-Stokes equations (1), reproduces available measurements $m_{n}^{R}=\mathcal{M}(\omega^{R}_{n})\in\mathbb{R}^{d_{m}}$ from a reference solution $\omega^{R}$ at discrete times $t_{n}=n\Delta t$ over the time horizon $t_{n}\in[0,T]$ for $n=0,\dots,N$. The problem is formulated as a variational minimisation of a cost function of the discrepancy between the estimated and true measurements,
@@
subject to the constraint that $\omega_{n}=f^{t_{n}}(\omega_{0})$ is the time-forward map of (1) from initial condition $\omega_{0}$. The required gradient of (2) with respect to $\omega_{0}$ can be computed using the discrete adjoint (wang2019discrete) or automatic differentiation (fan2025).
The measurement operator $\mathcal{M}$ is defined to be the coarse-graining operation $\mathcal{M}:\mathbb{R}^{N_{x}\times N_{y}}\to\mathbb{R}^{N_{x}/M\times N_{y}/M}$ which samples the high-resolution data at every $M^{\text{th}}$ gridpoint in both $x$- and $y$-directions. The temporal coarsening $\Delta t=M\delta t$ is set by the same coarsening factor, where $\delta t$ is the time-step of the numerical simulation. At $Re=\{40,100\}$, coarsening is set to $M=16$ and to $M=32$ at $Re=400$. The DA time horizon is $T\approx 0.6T_{L}$ where $T_{L}$ is the Lyapunov timescale at each $Re$.
@@
$\frac{\partial\omega^{\dagger}}{\partial\tau}+J(\psi,\omega^{\dagger})-\frac{1}{Re}\nabla^{2}\omega^{\dagger}+\psi^{\dagger}=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\omega}\,,\qquad\nabla^{2}\psi^{\dagger}-J(\omega,\omega^{\dagger})=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\psi}\,$, (3)
backwards in time, where $\tau:=T-t$, the streamfunction $\psi$ is related to the velocity components via $u=\partial_{y}\psi,v=-\partial_{x}\psi$, and $J(\psi,\omega)=\frac{\partial\psi}{\partial x}\frac{\partial\omega}{\partial y}-\frac{\partial\psi}{\partial y}\frac{\partial\omega}{\partial x}$. The adjoint field $\omega^{\dagger}$ at $t=0$ yields the variation of the cost function with respect to the initial condition $\omega_{0}$,
@@
which can be used to update the estimated $\omega_{0}$ in a direction that minimizes the cost by better reproducing the measurements. In lieu of explicit solution of the adjoint equations (3), we compute the gradient (4) by automatic differentiation. The cost function is then minimised using the Adam optimiser for a total of 500 optimisation steps at each $Re$. A sweep over initial step sizes (or learning rates) was performed, and $\alpha=0.2$ was selected.
Variational DA in state space requires a first guess of the initial flow field $\omega_{0}$. We will consider two initialisation approaches. The first is a bicubic interpolation of the measurements (InterpDA). The second is using a pre-trained super-resolution (SR) network $\mathcal{F}_{SR}:\mathbb{R}^{d_{m}}\to\mathbb{R}^{d_{\omega}}$, which maps from instantaneous measurements to instantaneous full resolution fields. This initialisation for data assimilation (SR-DA) was shown to yield significant improvement in the accuracy of the assimilation estimate by weyrauch2025. We therefore adopt their SR network for comparison. These assimilations in state space serve as a benchmark for the proposed latent-space assimilation.
3 Latent-space data assimilation
The classical variational-DA algorithm attempts to identify the optimal initial vorticity, or state-space representation of the flow, to reproduce the measurement. The main idea of our latent-space DA is to, instead, attempt to predict a latent-space representation $\eta\in\mathbb{R}^{d_{\eta}}$ (see the orange box in figure 1). Shifting the object of the optimisation to the latent space requires an evaluation of the gradient with respect to the new latent coordinates. The latent space is mapped to state space by a pre-trained decoder $\mathcal{F}_{\mathcal{D}}(\eta)=\omega\in\mathbb{R}^{d_{\omega}}$. The new DA cost function is then
@@ 3 Latent-space data assimilation
We note that no dynamical model in the latent space is required and all time marching is performed in state space, such that the estimated solution $f^{t}(\mathcal{F}_{\mathcal{D}}(\eta_{0}))$ for $t\in[0,T]$ exactly satisfies the Navier-Stokes equations (1). Furthermore, given a pre-existing adjoint solver for the flow in question, only the ability to compute vector-Jacobian products of the decoder is required to perform latent-space DA according to (6). As in state-space DA, the Adam optimiser is used to minimise (5), and the same number of optimisation steps were taken. A sweep over initial step sizes led to the choices $\alpha=\{5,0.2,0.01\}$ at $Re=\{40,100,400\}$.
3.1 The latent space
We adopt the latent space of the implicit rank-minimising autoencoder (IRMAE). Unlike standard autoencoders, IRMAE is distinguished by the addition of a series of fully-connected, linear layers in the bottleneck of the network, which drive down the intrinsic dimensionality of the latent representation (jing2020; Zeng_2024). The IRMAE architecture has demonstrated comparable performance to variational autoencoders for smooth interpolation in the latent space and generating new samples from random noise (jing2020). As such, the latent representation exhibits these favourable properties in addition to being approximately minimal rank.
@@ 3.1 The latent space
To train the networks at $Re=\{40,100,400\}$, datasets were generated by sampling long-time trajectories of the flow at every time unit, resulting in datasets with a total number of snapshots $N_{S}\approx\{6,10,10\}\times 10^{4}$, respectively. The networks are then trained to minimise the loss function
@@
where each $\omega^{j}$ is a full-field snapshot from the training dataset. Full details of the IRMAE architecture, training protocol and reconstruction accuracy of (7) are provided in the supplementary material. For each $Re$ considered, DA was performed on independent trajectories which were never seen during the training of the neural networks.
Variational DA in latent space requires an first guess of the initial latent representation $\eta_{0}$. This initialisation is obtained by using the aforementioned pre-trained SR network followed by the IRMAE encoder, i.e., $\eta_{0}=\mathcal{F}_{\mathcal{E}}\circ\mathcal{F}_{SR}(m^{R}_{0})$. Comparable performance can be achieved by training a fully-connected network to map directly from measurement space to the IRMAE latent space.
4 Results
4.1 Accuracy of the estimated fields
Figure 2: Relative error $\varepsilon$ of the estimated turbulent field by (blue) InterpDA, (orange) SR, (green) SR-DA and (red) LatentDA at (left) $Re=40$ with $M=16$, (middle) $Re=100$ with $M=16$ and (right) $Re=400$ with $M=32$ as a function of time normalised by the Lyapunov timescale. Bold lines: ensemble average of ten independent trajectories; Coloured regions: max/min values attained by the ensemble. Grey region marks the data-assimilation time horizon.
The accuracy of the flow trajectories is compared when the initial flow states are estimated from super-resolution (SR), super-resolution-initialised DA in state space (SR-DA), and DA in latent space (LatentDA). Standard state-space DA initialised with a bicubic interpolation (InterpDA) is presented as a benchmark. The evolution of the error,
@@ 4.1 Accuracy of the estimated fields
for each method at $Re=\{40,100,400\}$ is reported in figure 2, where $\omega^{*}_{0}$ and $\omega^{R}_{0}$ are the estimated and reference turbulent fields at $t=0$. The spatio-temporal coarsening factor was $M=16$ at both $Re=\{40,100\}$ and $M=32$ at $Re=400$, and DA was performed over the time horizon $T\approx 0.6T_{L}$ which is indicated by the grey shaded region in figure 2.
@@
Figure 3: Comparison of (a) reference $\omega^{R}_{0}$ at $t=0$ and the estimated field $\omega^{*}_{0}$ from (i-iv) InterpDA, SR, SR-DA, LatentDA at $Re=400$. (b) Contours of the out-of-plane vorticity, (c) enstrophy spectra $\Omega(k)$ as a function of wavenumber $k$. The black spectra and the black contours in (biv) denote the reference data.
@@
LatentDA (red) results in an improvement of approximately two orders of magnitude over the standard InterpDA approach (blue) and one order of magnitude over SR-DA (green) at $Re=\{40,100\}$. Even at $Re=400$, we still observe an improvement by more than one order of magnitude over the InterpDA approach, and a significant improvement over the best-performing approach in state space (SR-DA). Since the accuracy advantage of the estimated trajectories from LatentDA is retained throughout the observation horizon, and because the growth rate of $\varepsilon(t)$ after $t=T$ is consistent across all three DA approaches, accurate predictions from LatentDA can be made over much longer time intervals. For example, at $Re=100$ and using LatentDA, the prediction error is ${\sim}1\%$ at $2.5T_{L}$ (not shown), which is a four folds longer horizon than when using SR-DA.
@@
A comparison of the estimated $\omega^{*}_{0}$ fields at $Re=400$ using each approach is presented in figure 3. When the data is assimilated in state space (figure 3(bi)), $\omega^{*}_{0}$ exhibits unphysical high-wavenumber artefacts. In contrast, the estimated fields by SR (figure 3(bii)) is overly smooth, which is a symptom of the spectral bias of neural networks. Assimilating this field (SA-DA) introduces some high-frequency errors, but the enstrophy spectrum $\Omega(k)$ remains poorly representative of the true field. When the assimilation is performed in latent space, the predicted initial field is most accurate and $\Omega(k)$ is most consistent with the reference turbulent state. Similar trends were observed at the lower $Re$ considered.
@@
Figure 4: (a) The spectra $\Omega_{adj}$ of the (green) adjoint field in state space $\omega^{\dagger}$ and (red) latent adjoint field decoded to state space $\mathcal{F}_{\mathcal{D}}(\eta+\alpha\eta^{\dagger})-\mathcal{F}_{\mathcal{D}}(\eta)$. Contours of the (bi) state-space adjoint and (bii) decoded latent adjoint fields.

### v2

@@ Abstract
Starting from limited measurements of a turbulent flow, data assimilation (DA) attempts to estimate all the spatio-temporal scales of motion. Success is dependent on whether the system is observable from the measurements, or how much of the initial turbulent field is encoded in the available measurements. Adjoint-variational DA minimizes the discrepancy between the true and estimated measurements by optimizing the initial velocity or vorticity field (the ‘state space’). Here we propose to instead optimize in a lower-dimensional latent space which is learned by implicit rank minimizing autoencoders. Assimilating in latent space, rather than state space, redefines the observability of the measurements and identifies the physically meaningful perturbation directions which matter most for accurate prediction of the flow evolution. When observing coarse-grained measurements of two-dimensional Kolmogorov flow at moderate Reynolds numbers, the proposed latent-space DA approach estimates the full turbulent state with a relative error improvement of two orders of magnitude over the standard state-space DA approach. The small scales of the estimated turbulent field are predicted more faithfully with latent-space DA, greatly reducing erroneous small-scale velocities typically introduced by state-space DA. Furthermore, latent-space DA is demonstrated to be robust to noisy measurements at the range of Reynolds numbers considered. These findings demonstrate that the observability of the system from available data can be greatly improved when turbulent measurements are assimilated in the right space, or coordinates.
I Introduction
Assimilating experimental data into numerical simulations improves the fidelity of the simulations and enables nonintrusive access to all the scales of the estimated flow. However, the estimation of turbulence from limited measurements is a difficult ill-posed problem [26]. Turbulence presents challenges such as the chaotic nature of the forward and dual problems, the non-uniqueness of solutions consistent with the measurements, and the introduction of erroneous small-scale velocities that decay over the solution trajectory.
@@ I Introduction
The term latent space refers to a low-dimensional and interpretable representation of the turbulent field, whether by familiar modal decompositions [20], or by non-linear autoencoder transformations [1]. Physical insights of complex dynamical systems can be gleaned from these low-dimensional latent spaces [10]. Rank-minimizing autoencoders have successfully learned parsimonious representations of chaotic systems such as the Lorenz system, the Kuramoto-Sivashinsky equation and the lambda-omega reaction-diffusion system [27]. The latent spaces of such autoencoders have yielded insights on the nature of bursting events and the dynamical relevance of unstable periodic orbits in forced two-dimensional (2D) turbulence [4].
@@
In the context of turbulence estimation, data-driven methods have been used to directly super-resolve limited instantaneous observations to the full flow state [e.g. 9]. Another super-resolution study considered a time-history of scarce measurements and inferred the pressure field of forced isotropic turbulent flow [25]. Machine-learning and classical adjoint DA methods to estimate turbulence have been pursued mostly separately, with a few exceptions including the following examples: Du et al. [7] compared the estimation of wall turbulence using physics-informed neural networks and adjoint-variational techniques. Page [19] modified the training of super-resolution networks to incorporate a time-forward Navier-Stokes evolution in the output, and a comparison to future data. Most recently, Weyrauch et al. [24] used the output of this super-resolution network as an initial guess to adjoint-variational DA. While this last effort has interfaced data-driven methods and adjoint-variational DA, the two techniques were not fully integrated.
@@
Relatedly, some data-driven methods have been incorporated into statistical filtering approaches for data assimilation [18] and into adjoint-variational approaches to approximate the time forward operator [3], but neither of these cases exactly satisfy the governing dynamical equations. In this work, we combine a learned space that is discovered using data-driven methods with the physics constraints of adjoint-variational DA. This integration significantly enhances the observability of turbulence systems and their estimation across a wider range of scales, while still exactly satisfying the dynamical equations. In §II, we outline the flow configuration studied and summarize the standard variational DA procedure. In §III, we present the proposed latent-space DA procedure. In §IV, we compare and interpret the improved performance of latent-space DA. We conclude in §V.
II State-space data assimilation
@@
We consider Kolmogorov flow, which is monochromatically forced 2D turbulence on a square and doubly periodic domain [2]. The out-of-plane vorticity $\omega=\partial_{x}v-\partial_{y}u$ defines the flow state in state space and evolves according to
$\partial_{t}\omega+\bm{u}\cdot\boldsymbol{\nabla}{\omega}=\frac{1}{Re}\nabla^{2}\omega-k_{f}\cos{k_{f}y},$ (1)
where $\bm{u}=(u,v)$ is the velocity. In this non-dimensionalisation, the length scale $1/k^{*}=L^{*}/2\pi$ is the inverse of the fundamental wavenumber (asterisk denotes dimensional quantities). The time scale is $1/\sqrt{k^{*}\chi^{*}}$, where $\chi^{*}$ is the amplitude of the forcing in the momentum equation. The Reynolds number is therefore $Re\coloneq\sqrt{\chi^{*}/k^{*3}}/\nu$, where $\nu$ is the kinematic viscosity. The forcing wavenumber is set to $k_{f}=4$. Kolmogorov flow approaches an asymptotic regime beyond $Re\approx 50$ [5], and therefore we consider $Re=\{40,100,400\}$ in this work. The vorticity-velocity Navier-Stokes equations (1) are solved using the pseudospectral version of the JAX-CFD solver [16, 6], allowing for the efficient computation of gradients of the time-forward map of (1) using automatic differentiation. The computational grid is set to $N_{x}\times N_{y}=128^{2}$ for $Re=\{40,100\}$ and $512^{2}$ for $Re=400$.
@@
Figure 1: Schematic of latent-space data assimilation. The latent representation $\eta$ is mapped to state space $\omega$ by the decoder $\mathcal{F}_{\mathcal{D}}$, where the adjoint field $\omega^{\dagger}$ is computed. The gray arrow closes the standard, state-space ‘loop’. In latent-space assimilation, the latent state is updated by the transformed adjoint field $\eta^{\dagger}$. Sensor resolution is indicated by the lattice of black dots in the lower left corner of $\omega$.
@@
Our objective is to estimate the flow state $\omega^{*}_{0}$ which, when evolved using the Navier-Stokes equations (1), reproduces available measurements $m_{n}^{R}=\mathcal{M}(\omega^{R}_{n})\in\mathbb{R}^{d_{m}}$ from a reference solution $\omega^{R}$ at discrete times $t_{n}=n\Delta t$ over the time horizon $t_{n}\in[0,T]$ for $n=0,\dots,N$. The problem is formulated as a variational minimization of a cost function of the discrepancy between the estimated and true measurements,
@@
subject to the constraint that $\omega_{n}=f^{t_{n}}(\omega_{0})$ is the time-forward map of (1) from initial condition $\omega_{0}$. The required gradient of (2) with respect to $\omega_{0}$ can be computed using the discrete adjoint [21] or automatic differentiation [8].
The measurement operator $\mathcal{M}$ is defined to be the coarse-graining operation $\mathcal{M}:\mathbb{R}^{N_{x}\times N_{y}}\to\mathbb{R}^{N_{x}/M\times N_{y}/M}$ which samples the high-resolution data at every $M^{\text{th}}$ gridpoint in both $x$- and $y$-directions. The temporal coarsening $\Delta t=M\delta t$ is set by the same coarsening factor, where $\delta t$ is the time-step of the numerical simulation. At $Re=\{40,100\}$, coarsening is set to $M=16$ and to $M=64$ at $Re=400$, resulting in $8^{2}$ spatial measurements for each $Re$ case. The DA time horizon is $T\approx 0.6T_{L}$ where $T_{L}$ is the Lyapunov timescale at each $Re$.
@@
$\frac{\partial\omega^{\dagger}}{\partial\tau}+J(\psi,\omega^{\dagger})-\frac{1}{Re}\nabla^{2}\omega^{\dagger}+\psi^{\dagger}$ $=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\omega}\,$, (3a)
$\nabla^{2}\psi^{\dagger}-J(\omega,\omega^{\dagger})$ $=\frac{\mathcal{D}\mathcal{J}}{\mathcal{D}\psi}\,$, (3b)
backwards in time, where $\tau:=T-t$, as derived in Appendix A. The streamfunction $\psi$ is related to the velocity components via $u=\partial_{y}\psi,v=-\partial_{x}\psi$, and $J(\psi,\omega)=\frac{\partial\psi}{\partial x}\frac{\partial\omega}{\partial y}-\frac{\partial\psi}{\partial y}\frac{\partial\omega}{\partial x}$. The adjoint field $\omega^{\dagger}$ at $t=0$ yields the variation of the cost function with respect to the initial condition $\omega_{0}$,
@@
which can be used to update the estimated $\omega_{0}$ in a direction that minimizes the cost by better reproducing the measurements. In lieu of explicit solution of the adjoint equations (3), we compute the gradient (4) by automatic differentiation. The cost function is then minimized using the Adam optimizer [15] for a total of 500 optimization steps at each $Re$. A sweep over initial step sizes (or learning rates) was performed, such that $\alpha=0.2$ was selected as the initial step size at each $Re$.
Variational DA in state space requires a first guess of the initial flow field $\omega_{0}$. We will consider two initialization approaches. The first is a bicubic interpolation of the measurements (InterpDA). The second is using a pre-trained super-resolution (SR) network $\mathcal{F}_{SR}:\mathbb{R}^{d_{m}}\to\mathbb{R}^{d_{\omega}}$, which maps from instantaneous measurements to instantaneous full resolution fields. This initialization for data assimilation (SR-DA) was shown to yield significant improvement in the accuracy of the assimilation estimate by Weyrauch et al. [24]. The same network as in Page [19] is adopted here, which is trained on the standard $L_{2}$ loss between the full resolution snapshot and the super-resolved snapshot. These SR networks were trained with the Adam optimizer with an initial learning rate of $10^{-4}$ and a batch size of 32, on data sets of 5000 turbulent snapshots at each $Re$ for a total of 250 epochs. Full details on the data generation is given in Appendix C. These assimilations in state space serve as a benchmark for the proposed latent-space assimilation.
III Latent-space data assimilation
The classical variational-DA algorithm attempts to identify the optimal initial vorticity, or state-space representation of the flow, to reproduce the measurement. The main idea of our latent-space DA is to, instead, attempt to predict a latent-space representation $\eta\in\mathbb{R}^{d_{\eta}}$ (see the orange box in figure 1). Shifting the object of the optimization to the latent space requires an evaluation of the gradient with respect to the new latent coordinates. The latent space is mapped to state space by a pre-trained decoder $\mathcal{F}_{\mathcal{D}}(\eta)=\omega\in\mathbb{R}^{d_{\omega}}$. The new DA cost function is then
@@ III Latent-space data assimilation
We note that no dynamical model in the latent space is required and all time marching is performed in state space, such that the estimated solution $f^{t}(\mathcal{F}_{\mathcal{D}}(\eta_{0}))$ for $t\in[0,T]$ exactly satisfies the Navier-Stokes equations (1). Furthermore, given a pre-existing adjoint solver for the flow in question, only the ability to compute vector-Jacobian products of the decoder is required to perform latent-space DA according to (6). As in state-space DA, the Adam optimizer is used to minimize (5), and the same number of optimization steps were taken. A sweep over initial step sizes led to the choices $\alpha=\{5,0.2,0.02\}$ at $Re=\{40,100,400\}$.
III.1 The latent space
We adopt the latent space of the implicit rank-minimizing autoencoder (IRMAE). Unlike standard autoencoders, IRMAE is distinguished by the addition of a series of fully-connected, linear layers in the bottleneck of the network, which drive down the intrinsic dimensionality of the latent representation [14, 27]. The IRMAE architecture has demonstrated comparable performance to variational autoencoders for smooth interpolation in the latent space and generating new samples from random noise [14]. As such, the latent representation exhibits these favorable properties in addition to being approximately minimal rank.
@@ III.1 The latent space
To train the networks at $Re=\{40,100,400\}$, datasets were generated by sampling long-time trajectories of the flow at every time unit, resulting in datasets with a total number of snapshots $N_{S}\approx\{6,10,10\}\times 10^{4}$, respectively. Initial transient periods of 50 time units were discarded from the trajectories, such that the training data is sampled from the turbulent attractor. The networks are then trained to minimize the loss function
@@
where each $\omega^{j}$ is a full-field snapshot from the training dataset. Full details of the IRMAE architecture and training protocol and reconstruction accuracy of (7) are provided in Appendices B and C, respectively. For each $Re$ considered, DA was performed on independent trajectories which were never seen during the training of the neural networks.
Variational DA in latent space requires an first guess of the initial latent representation $\eta_{0}$. This initialization is obtained by using the aforementioned pre-trained SR network followed by the IRMAE encoder, i.e., $\eta_{0}=\mathcal{F}_{\mathcal{E}}\circ\mathcal{F}_{SR}(m^{R}_{0})$. Comparable performance can be achieved by training a fully-connected network to map directly from measurement space to the IRMAE latent space.
IV Results
IV.1 Accuracy of the estimated fields
Figure 2: Relative error $\varepsilon$ of the estimated turbulent field by (blue) InterpDA, (orange) SR, (green) SR-DA and (red) LatentDA at (left) $Re=40$, (middle) $Re=100$ and (right) $Re=400$ as a function of time normalized by the Lyapunov timescale. Bold colored lines: ensemble mean of ten independent trajectories. Gray region marks the data-assimilation time horizon.
The accuracy of the flow trajectories is compared when the initial flow states are estimated from a single inference of the super-resolution network (SR), super-resolution-initialized DA in state space (SR-DA), and DA in latent space (LatentDA). Standard state-space DA initialized with a bicubic interpolation (InterpDA) is presented as a benchmark. The evolution of the error,
@@ IV.1 Accuracy of the estimated fields
for each method at $Re=\{40,100,400\}$ is reported in figure 2, where $\omega^{*}_{0}$ and $\omega^{R}_{0}$ are the estimated and reference turbulent fields at $t=0$. The spatio-temporal coarsening factor was $M=16$ at both $Re=\{40,100\}$ and $M=64$ at $Re=400$, and DA was performed over the time horizon $T\approx 0.6T_{L}$ which is indicated by the gray shaded region in figure 2.
@@
Figure 3: Comparison of the estimated field $\omega^{*}_{0}$ from (i-iv) InterpDA, SR, SR-DA and LatentDA (highlighted by gray box) with (c) reference $\omega^{R}_{0}$ at $t=0$ at $Re=400$. The lattice of black dots in (c) indicates the sensor resolution. (b) Enstrophy spectra $\Omega(k)$ as a function of wavenumber $k$. The black spectra denote the reference data.
ReRe InterpDA SR SR-DA LatentDA
 40    2.5 $\pm$  1.7   22.4 $\pm$ 6.5   0.4 $\pm$ 0.8   0.06 $\pm$ 0.04
100   15.0 $\pm$ 19.2   60.7 $\pm$ 4.4   1.6 $\pm$ 0.3   0.28 $\pm$ 0.04
400   51.6 $\pm$ 25.8   65.5 $\pm$ 6.8   7.5 $\pm$ 4.2   1.70 $\pm$ 1.10
Table 1: Mean and standard deviation of the percent relative error ($\varepsilon\times 100$) of the estimated turbulent field at the DA time horizon $t=T$, over an ensemble of ten independent trajectories.
@@
LatentDA (red) results in an improvement of approximately two orders of magnitude over the standard InterpDA approach (blue) and one order of magnitude over SR-DA (green) at $Re=\{40,100\}$. Even at $Re=400$, we still observe an improvement by more than one order of magnitude over the InterpDA approach, and a significant reduction of ${\sim}50\%$ error over the best-performing approach in state space (SR-DA). Furthermore, table 1 reveals that the standard deviation of the percent relative error at the DA time horizon $T$ using LatentDA is significantly smaller than SR-DA at $Re=400$. Since the accuracy advantage of the estimated trajectories from LatentDA is retained throughout the observation horizon, and because the growth rate of $\varepsilon(t)$ after $t=T$ is consistent across all three DA approaches, accurate predictions from LatentDA can be made over much longer time intervals. For example, at $Re=100$ and using LatentDA, the prediction error is ${\sim}1\%$ at $2.5T_{L}$ (not shown), which is a four folds longer horizon than when using SR-DA.
@@
A comparison of the estimated $\omega^{*}_{0}$ fields at $Re=400$ using each approach is presented in figure 3. When the data is assimilated in state space (figure 3i), $\omega^{*}_{0}$ exhibits unphysical high-wavenumber artifacts and does not at all resemble the reference data (figure 3c). In contrast, the estimated field by SR (figure 3ii) is overly smooth, which is a symptom of the spectral bias of neural networks. State-space DA initialized with this field (SR-DA) introduces some high-frequency errors (figure 3iii), but the enstrophy spectrum $\Omega(k)$ remains poorly representative of the true field. When the assimilation is performed in latent space (figure 3iv, highlighted by the gray box), the predicted initial field is most accurate and $\Omega(k)$ is most consistent with the reference turbulent state. Similar trends were observed at the lower $Re$ considered.
@@
Figure 4: (a) The spectra $\Omega_{adj}$ of the initial normalized adjoint field in state space $\omega^{\dagger}$ for (blue) InterpDA, (green) SR-DA and (red) LatentDA adjoint field decoded to state space $\mathcal{F}_{\mathcal{D}}(\eta+\alpha\eta^{\dagger})-\mathcal{F}_{\mathcal{D}}(\eta)$ at $Re=400$. Contours of the state-space adjoint for InterpDA (bi), SR-DA (bii) and (biii) decoded LatentDA adjoint fields.